The half-life of cesium-137 is 30 years. Suppose we have a 89 g sample. Find a function that models the mass m remaining after t years.
Answer #1

-at No= 89gm We know N= No e -2 tl/2 No - No e 그 1 = 0.69314716 + 4,3909g 0.693 14728 30 ti Let N= m (mars remaining after t